I borrow £5000 and agree to pay back £6000 after 18 months. Whatever is the annual rate of compound interest?



What is asked:



What are given:



What is the answer:

1
Expert's answer
2021-11-19T11:44:22-0500

Variable Asked:

Annual Rate of Compound Interest=i

Borrowed amount (p)=5000

Repayment amount(A)=6000

Period (years)t="\frac{months}{12}"

"=\frac{18}{12}=1.5"

"A=p(1+i)^t\\6000=5000×(1+i)^{1.5}\\1+i=(\frac{6000}{5000})^{\frac{1}{1.5}}\\
1+i=1.129243\\i=1.129243−1\\i=0.129243\\i=12.9243\%"

So annual rate compounded is 12.9243%